OpenStudy (anonymous):

(sqrt3x-1)^2= (2x)^2 help solve please

OpenStudy (mertsj):

\[(\sqrt{3x-1})^{2}=3x-1\]

OpenStudy (vishal_kothari):

imaginary roots

OpenStudy (mertsj):

So 3x-1 = 4x^2 4x^2-3x+1=0

OpenStudy (mertsj):

Now use the quadratic formula.

OpenStudy (anonymous):

No, vishal, this doesn't require imaginary numbers because we aren't taking the root of a negative number

OpenStudy (mertsj):

Yes we are. since 9-16 is negative.

OpenStudy (anonymous):

@mertsj you are right...

OpenStudy (mertsj):